OpenStudy (anonymous):

Hi everyone, I'm currently studying for my calculus test tommorow and I have some problems which I have no clue how to solve. 1) Lim (cos 2x)^(csc^2 2x) x-> 0 Any help at all is very much appreciated, as I truly have no idea how to get started this excercise. I'm sorry for not saying this earlier but L'Hopitals rule is not allowed.

OpenStudy (anonymous):

this looks like \(1^{\infty}\) so you can use l'hopital after you take the log

OpenStudy (anonymous):

when you see an exponential function like this, you usually have to take the log first and work with that one either that, or rewrite as \[e^{\csc^2(2x)\ln(\cos(2x)}\]

OpenStudy (anonymous):

i would take the log, get \[\csc^2(2x)\ln(\cos(2x))\] then rewrite as \[\frac{\ln(\cos(2x))}{\sin^2(2x)}\] to put it in the form \(\frac{0}{0}\) now you can use l'hopital

OpenStudy (anonymous):

when you get your answer to this, make sure to remember to raise \(e\) to that power, as you took the log as the first step

OpenStudy (anonymous):

Thank you very much for your fast answer but unfortunately we are not allowed to use L'Hoptials rule yet. I'm sorry I didn't specify that, is there another way to solve this problem?

OpenStudy (anonymous):

i dont know whether it help take the whole expression = y and take log on both sides try to simplify

OpenStudy (anonymous):

That looks like it could work. I'll try that.

OpenStudy (anonymous):

But isn't that for solving derivatives?

OpenStudy (anonymous):

Ok so I got this one in the form (1-2sin^2 x)^(1/2sin^2 2x) but i don't know how to go on from there. I know i'm almost there but i don't know how to get rid of the 2x above, I know it's silly, but i just don't see it.

OpenStudy (anonymous):

1-2sin^2x = cos2x